OpenStudy (anonymous):

Simplify: (7-3) ÷ 2 + 4 × 6

OpenStudy (yanasidlinskiy):

\(\huge\cal\color{Lime}{Welcome~to~OpenStudy!!!!!}\) Remove the parenthases by subtracting 7-3. On the other side, multiply the 6 and the 4. Then add that answer to the 2. Finally, divide.

OpenStudy (anonymous):

thx

OpenStudy (yanasidlinskiy):

Your Welcome!!:) ~Stay in touch with OpenStudy!:)~

OpenStudy (anonymous):

@iGreen i need help

OpenStudy (anonymous):

skylar said to help

OpenStudy (yanasidlinskiy):

What?

OpenStudy (anonymous):

juss making sure

OpenStudy (igreen):

PEMDAS Okay, first you simplify the parenthesis. 7 - 3 = 4. 4 ÷ 2 + 4 × 6 Then you divide. 4 ÷ 2 = 2 2 + 4 × 6 Then multiply 4 × 6 = 24 2 + 24 Then add 26